oslo company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 to 1,500): sales $85,00, variable expensed 59,500, contribution margin 25,500 fixed expensed 20,400 net operating income $51,100. how many units must be sold to achieve a target profit of $15,300
Answer #1

Contribution margin per unit = Total contribution margin/Number of units sold

= 25,500/1,000

= $25.5

Number of units to be sold to earn target profit = (Fixed cost + Target profit)/Contribution margin per unit

= (20,400 + 15,300)/25.5

= 35,700/25.5

= 1,400 units

Units must be sold to achieve a target profit of $15,300 = 1,400 units

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.